In your sock drawer , there are 30 black socks, 12 white socks, and 20 brown socks. To the nearest hundredth, what is the probability of randomly selecting two white socks without replacement
Answer by lynnlo(4176)

p(30)p(12)p(20)
p(2)=(12/62)*(11/61)
p(2)=132/3782
p(2)=0.03
multiply================12*11=132
multiply 62*61=3782
divide 132 by 3782==0.03